OpenStudy (anonymous):

How would you simplify 3x^3 + 9z +2x^3 +5z +6x^2 by combining like terms?

OpenStudy (anonymous):

Would the answer be 11x^3+14z ?

OpenStudy (anonymous):

\[(3x^3+2x^3)+(6x^2) +(9z+5z)\] every thing in a bracket is a "like term" try again

OpenStudy (anonymous):

Well I guessed pretty much . I have no idea how to do this .Could you possibly explain how to ..?

OpenStudy (anonymous):

so first you need to understand what a like term is..... 3, 4, 5, 6 are all CONSTANTS (like terms) 3x 5x 6x 7x are all like terms (only have one x value) so basically they have to belong to the same family to be a like term... can you give me another example of some like terms?

OpenStudy (anonymous):

6x ,8x,10x

OpenStudy (anonymous):

how about \[x^3\] and \[x^4\] are they like terms?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

you sound unsure, \[x^3=x \times x \times x\] \[x^4=x \times x \times x \times x\] do you still think they are like terms?

OpenStudy (anonymous):

.....

OpenStudy (anonymous):

5x^3+14z+6x^2